# LLM-Driven Corrective Robot Operation Code Generation with Static Text-Based Simulation

### v2

@@ IV-D Ablation Study on LLM-Simulator Design
We conduct an ablation study on the design of our LLM-simulator. Specifically, we investigate how different components of the LLM’s system prompt affect the overall system performance. As specified in Section III-D, the role defines the LLM to be a simulator and the APIs provide definitions of robot action APIs, they serve as the essential component to ensure the LLM performs the text-based simulation. Thus, they are retained in the ablation study. For the remaining components, we remove polices, examples, and both polices and examples to measure their impact on the overall system performance. The results in Table V show that removing either polices or examples leads to degradation in performance, while removing both yields the largest decline. Furthermore, removing examples produces a larger negative impact than removing polices, which is consistent with prior evidence that LLMs exhibit strong few-shot learning capabilities [llm-fewshot]. The results of the ablation study highlight that both instructing the LLM with code execution rules (polices) and providing demonstrations (examples) are essential for enhancing the performance of the LLM-simulator, and their combination yields the best overall performance.
@@
Ours w/o w/o w/o polices
polices examples & examples
@@ V-A1 UAV
We deploy our method on the Holybro X500 V2 quadcopter equipped with a Pixhawk 6X flight controller running PX4 firmware. For the task set, we select two tasks from each complexity level of the Advanced task set, resulting in a total of 8 tasks for deployment. In the deployment, the user types the task description into a ground station computer that runs our method to generate the UAV operation code (the computer is connected to the Internet for accessing the OpenAI API). The generated code is then transmitted to the flight controller via MAVSDK [mavsdk] for task execution. During the flight, the UAV’s state information is collected through MAVSDK for performance measurement.
@@ V-A2 Ground Vehicle
We further evaluate the method on a ROSMASTER X3 ground vehicle controlled by ROS [ROS]. We then design 8 tasks for the ground vehicle to form patterns when driving on the ground. In the deployment, the user accesses the onboard computer remotely and types in the task descriptions, then the onboard computer (connected to the Internet for accessing the OpenAI API) runs our method to generate the ground vehicle operation code and then executes the code to control the vehicle’s movement.
@@ VI-A LLM-Simulator System Prompt
Polices: Important drone coordinate directional information and action polices: 1. The horizontal axes are Y and X, the vertical axis is Z. 2. When rotating the drone, turning right or clockwise means positive, the yaw angle should increase. 3. aw.fly_to([x, y, z]) function uses NED coordinate system (world coordinates), positive X axis is North/forward, positive Y axis is East/right, positive Z axis is Down. When flying up, the Z value should decrease. When flying down, the Z value should increase. 4. The drone is initialized facing north (Yaw = 0 degrees). 5. Map Yaw angle degree from -180 to 180, for example: map 270 to -90.
@@ VI-B Simulator Configurations
In our experiments, we employ two widely used simulators: AirSim [AirSim] and Gazebo [gazebo].
@@ VI-B1 AirSim
Developed by Microsoft, AirSim provides a high-fidelity virtual environment for research in autonomous systems such as UAVs, self-driving cars, and ground robots. Built on Unreal Engine [unreal], it supports realistic rendering of robot dynamics. In our experiments, we use the pre-built “block” [Blocks] environment, which provides a clear and open airspace suitable for UAV testing.
@@ VI-B2 Gazebo
Gazebo [gazebo] models physical dynamics such as gravity, friction, and contact forces, making it applicable to both aerial and ground robots. It integrates seamlessly with the ROS operating system [ROS]. For our experiments, we use the PX4 [PX4] software-in-the-loop (SITL) setup to control a simulated quadcopter within a virtual world environment of Gazebo.

### v3

@@ IV-D Ablation Study on LLM-Simulator Design
We conduct an ablation study on the design of our LLM-simulator. Specifically, we investigate how different components of the LLM’s system prompt affect the overall system performance. As specified in Section III-D, the role defines the LLM to be a simulator, and the APIs provide definitions of robot action APIs, they serve as the essential component to ensure the LLM performs the text-based simulation. Thus, they are retained in the ablation study. For the remaining components, we remove policies, examples, and both policies and examples to measure their impact on the overall system performance. The results in Table V show that removing either policies or examples leads to degradation in performance, while removing both yields the largest decline. Furthermore, removing examples produces a larger negative impact than removing policies, which is consistent with prior evidence that LLMs exhibit strong few-shot learning capabilities [4]. The results of the ablation study highlight that both instructing the LLM with code execution rules (policies) and providing demonstrations (examples) are essential for enhancing the performance of the LLM-simulator, and their combination yields the best overall performance.
@@
Ours w/o w/o w/o policies
policies examples & examples
@@ V-A1 UAV
We deploy our method on the Holybro X500 V2 quadcopter equipped with a Pixhawk 6X flight controller running PX4 firmware. For the task set, we select two tasks from each complexity level of the Advanced task set, resulting in a total of 8 tasks for deployment. In the deployment, the user types the task description into a ground station computer that runs our method to generate the UAV operation code (the computer is connected to the Internet for accessing the OpenAI API). The generated code is then transmitted to the flight controller via MAVSDK [18] for task execution. During the flight, the UAV’s state information is collected through MAVSDK for performance measurement.
@@ V-A2 Ground Vehicle
We further evaluate the method on a ROSMASTER X3 ground vehicle controlled by ROS [25]. We then design 8 tasks for the ground vehicle to form patterns when driving on the ground. In the deployment, the user accesses the onboard computer remotely and types in the task descriptions, then the onboard computer (connected to the Internet for accessing the OpenAI API) runs our method to generate the ground vehicle operation code and then executes the code to control the vehicle’s movement.
@@ VI-A LLM-Simulator System Prompt
Policies: Important drone coordinate directional information and action policies: 1. The horizontal axes are Y and X, the vertical axis is Z. 2. When rotating the drone, turning right or clockwise means positive, the yaw angle should increase. 3. aw.fly_to([x, y, z]) function uses NED coordinate system (world coordinates), positive X axis is North/forward, positive Y axis is East/right, positive Z axis is Down. When flying up, the Z value should decrease. When flying down, the Z value should increase. 4. The drone is initialized facing north (Yaw = 0 degrees). 5. Map Yaw angle degree from -180 to 180, for example: map 270 to -90.
@@ VI-B Simulator Configurations
In our experiments, we employ two widely used simulators: AirSim [27] and Gazebo [13].
@@ VI-B1 AirSim
Developed by Microsoft, AirSim provides a high-fidelity virtual environment for research in autonomous systems such as UAVs, self-driving cars, and ground robots. Built on Unreal Engine [8], it supports realistic rendering of robot dynamics. In our experiments, we use the pre-built “block” [20] environment, which provides a clear and open airspace suitable for UAV testing.
@@ VI-B2 Gazebo
Gazebo [13] models physical dynamics such as gravity, friction, and contact forces, making it applicable to both aerial and ground robots. It integrates seamlessly with the ROS operating system [25]. For our experiments, we use the PX4 [19] software-in-the-loop (SITL) setup to control a simulated quadcopter within a virtual world environment of Gazebo.